# Boundary Control and Calderón type Inverse Problems in Non-local heat equation

### v2

@@ 1.2. Well posedness of the solution
[…]
@@
Let $P(t,x,y)$ be the fractional heat semigroup corresponding to the operator $\partial_{t}+(-\Delta)^{a}+q$ with homogeneous Dirichlet boundary condition and $\delta_{x}$(Dirac delta) as initial condition. Then the quantity ${\lim\limits_{\begin{subarray}{c}y\to w\\ w\in\partial\Omega\end{subarray}}\frac{P(t,x,y)}{d^{a-1}(y)}}$ is a well defined quatity [see [CGCV22]]. Well defineness of that particular quantity comes from existence of the same limit function for the corresponding elliptic Green function with homogeneous Dirichlet boundary condition [see [AGCV23]]. The solution $u_{f}$ corresponding to (1) or $u_{F}$ corresponding to (2) can be uniquely expressed as
[…]
$u_{F}(t,x)=\int_{0}^{t}\int_{\partial\Omega}\lim\limits_{\begin{subarray}{c}y\to w\\ w\in\partial\Omega\end{subarray}}\frac{S(t-s,x,y)}{d^{a-1}(y)}h(s,w)$
@@
where $\lambda_{1}<\lambda_{2}\leq\dots\leq\lambda_{n}\leq\dots+\infty$ are the eigenvalues and $\phi_{n}$’s are the eigenfunctions of the following Dirichlet valued problem
(6) $\left\{\begin{array}[]{ll}((-\Delta)^{a}+q)\phi_{n}=\lambda_{n}\phi_{n}&\mbox{in}\ \Omega\\ \phi_{n}=0&\mbox{in}\ \mathbb{R}^{N}\setminus\Omega.\end{array}\right.$
@@
  • ((((a))))
    For eact $t_{0}>0$
    (7) $\sup\limits_{t\geq t_{0}}\|u_{f}(t,\cdot)\|_{C_{c}^{a}(\mathbb{R}^{N})}\leq C_{1}(t_{0})\|f\|_{\mathrm{L}^{2}(\Omega)}.$
  • ((((b))))
    For eact $t_{0}>0$
@@
for any $\epsilon\in(0,a)$. Here $C_{1}(t_{0})$,$C_{1}^{j}(t_{0})$ are positive constants. Authors used various eigenfunction estimates corresponding to (6) [see [Gru15b], [Gru15a], [AGCV23]]. The following interior regularity results can be found in [FRRO16].
@@
Let $u_{f}\in\mathrm{L}^{2}\left((0,T);H^{a}(\Omega)\right)$ be the solution of (1) with $f\in\mathrm{L}^{2}(\Omega)$. Then the following holds
@@
    For $t>0$, $u(t,\cdot)\in C^{a}(\mathbb{R}^{N})$ and, for every $\beta\in[a,1+2a)$, $u(t,\cdot)$ is of class $C^{\beta}(\Omega)$ and
    (10) $[u(t,\cdot)]_{C^{\beta}(\Omega_{\delta})}\leq C\delta^{a-\beta},\forall\delta\in(0,1).$
@@
    The function ${\frac{u(t,\cdot)}{d^{a}}}\Big{|}_{\Omega}$ can be continuously extended to $\overline{\Omega}$. Moreover, there exists $\alpha\in(0,1)$ such that ${\frac{u}{d^{a}}\in C^{\alpha}(\overline{\Omega})}$. In addition, for $\beta\in[\alpha,a+\alpha]$, the following estimate holds
    (11) $\left[\frac{u}{d^{a}}\right]_{C^{\beta}(\Omega_{\delta})}\leq C\delta^{a-\beta},\forall\delta\in(0,1).$
Here $\Omega_{\delta}:=\{x\in\Omega:d(x,\partial\Omega)\geq\delta\}$ and $[\cdot]_{C^{\beta}(\Omega_{\delta})}$ denotes the Hölder seminorm:
$\left[\frac{u(t,x)}{d^{a}}\right]_{C^{\beta}(\Omega_{\delta})}=\sup\limits_{x,y\in C^{\beta}(\Omega_{\delta}),x\neq y}\frac{|D^{\lfloor\beta\rfloor}u(t,x)-D^{\lfloor\beta\rfloor}u(t,y)|}{|x-y|^{\beta-\lfloor\beta\rfloor}}$
@@
(12) $\int_{\Omega}(x\cdot\nabla u)(-\Delta)^{a}u\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}$ $u(-\Delta)^{a}u\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{u}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma,$
@@ 1.3. $\mu$ transmission spaces
All the notations and definition here are well explained in [Gru14], [Gru15a], [Gru16b], [Gru16a], [Gru18], [Gru20]. We only recall some important definitions and concepts here.
@@
$\mathcal{E}_{\mu}\left(\overline{\Omega}\right):=e^{+}\left\{u(x)=\tilde{d}^{\mu}v(x):v\in C^{\infty}\left(\overline{\Omega}\right)\right\},$
for Re$(\mu)>-1$. For general $\mu$, we refer readers to [Gru15a], [See71]. Next we define pseudodifferential operator ($\psi$do) $\mathcal{P}$ on $\mathbb{R}^{N}$, which is defined from a symbol $p(x,\xi)$ on ${\mathbb{R}^{N}\times\mathbb{R}^{N}}$ in the following way:
@@
$S^{m}_{1,0}:=\left\{\begin{aligned} p(x,\xi)\in C^{\infty}\left(\mathbb{R}^{N}\times\mathbb{R}^{N}\right):&|\partial_{x}^{\beta}\partial_{x}^{\alpha}p(x,\xi)|\sim O\left(\langle\xi\rangle^{m-|\alpha|}\right),\\ &\forall\alpha,\beta\in\mathbb{N}\cup\{0\},\ \mbox{for some}\ m\in\mathbb{C},\end{aligned}\right.$
where ${\langle\xi\rangle:=\left(1+\|\xi\|^{2}_{2}\right)^{\frac{1}{2}}}$ and $O$ is the big-O function captures the order. The exponent $m$ is called the order of $\mathcal{P}($and $p)$. Note that the operator $\left((-\Delta)^{a}+q\right)$ is a $\psi$do operator of order 2a. Depending on the context, the operator $\mathcal{P}_{+}=r^{+}\mathcal{P}e^{+}$ denotes the truncation of $\mathcal{P}$ to +$\mathbb{R}^{N}_{+}$ or to $\Omega$. We define the following $\mathrm{L}^{2}$ spaces:
@@
The space ${{\mathcal{E}_{\mu}\left(\overline{\Omega}\right)}}$ is dense in ${H^{\mu(s)}\left(\overline{\Omega}\right)}$ i.e. [see [Hor63], [Hör66], [Gru08]]
@@
Let $f\in C_{c}^{\infty}(\Omega)$, then the solution $u_{f}$ to (1) has the following regularity:
$u_{f}\in\mathrm{L}^{2}\left((0,T);H^{a(2a)}(\overline{\Omega})\right).$
It follows from [Gru23]. The function $v=u_{f}-f$ satisfies
$\partial_{t}v+(-\Delta)^{a}v+qv=&(-\Delta)^{a}f+qf\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ v=&0\qquad\qquad\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ v(0,\cdot)=&0\qquad\qquad\qquad\qquad\qquad\mbox{in}\ \Omega.\end{aligned}\right.$
As $f\in C_{c}^{\infty}(\Omega)$, we have $(-\Delta)^{a}f+qf\in\mathrm{L}^{2}((0,T)\times\Omega)$. Hence $v\in\mathrm{L}^{2}\left((0,T);H^{a(2a)}(\overline{\Omega})\right)$ [Gru23]. Furthermore, as $f\in\mathcal{E}_{a}(\overline{\Omega})$, we have
$u_{f}\in\mathrm{L}^{2}\left((0,T);H^{a(2a)}(\overline{\Omega})\right).$
@@
All the above results are also true for the solution of the adjoint system corresponding to (1), given by:
(13) $-\partial_{t}u_{f^{*}}+(-\Delta)^{a}u_{f^{*}}+qu_{f^{*}}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{f^{*}}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ u_{f^{*}}(T,\cdot)=&f^{*}\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $f^{*}\in\mathrm{L}^{2}(\Omega)$. It just comes from change of variable $t\to T-t$. We like to emphasize that Pohozaev identity (12) holds for $u_{f^{*}}$ also.

### v3

@@ 1.2. Wellposedness of the solution
$\mathrm{L}^{1}\left(\Omega,d^{s}\right):=\left\{f:\int_{\Omega}fd^{s}dx<+\infty\right\}.$
@@
Let $P(t,x,y)$ be the fractional heat semigroup corresponding to the operator $\partial_{t}+(-\Delta)^{a}+q$ with homogeneous Dirichlet boundary condition and $\delta_{x}$(Dirac delta) as initial condition. Then the quantity ${\lim\limits_{\begin{subarray}{c}y\to w\\ w\in\partial\Omega\end{subarray}}\frac{P(t,x,y)}{d^{a-1}(y)}}$ is a well defined quatity [see [CGCV22]]. Well defineness of that particular quantity comes from the existence of the same limit function for the corresponding elliptic Green function with homogeneous Dirichlet boundary condition [see [AGCV23]]. The solution $u_{f}$ corresponding to (1) or $u_{F}$ corresponding to (2) can be uniquely expressed as
$u_{f}(t,x)=\int_{\Omega}P(t,x,y)f(y)dy$
$u_{F}(t,x)=\int_{0}^{t}\int_{\partial\Omega}\lim\limits_{\begin{subarray}{c}y\to w\\ w\in\partial\Omega\end{subarray}}\frac{S(t-s,x,y)}{d^{a-1}(y)}F(s,w)\,\rm{d}w\,\rm{d}s.$
@@
where $\lambda_{1}<\lambda_{2}\leq\dots\leq\lambda_{n}\leq\dots+\infty$ are the eigenvalues and $\phi_{n}$’s are the normalized eigenfunctions of the following homogeneous Dirichlet problem
(6) $\left\{\begin{array}[]{ll}((-\Delta)^{a}+q)\phi_{n}=\lambda_{n}\phi_{n}\qquad\qquad&\mbox{in}\ \Omega\\ \phi_{n}=0&\mbox{in}\ \mathbb{R}^{N}\setminus\Omega.\end{array}\right$.
@@
  • ((a))
    for each $t_{0}>0$
    (7) $\sup\limits_{t\geq t_{0}}\|u_{f}(t,\cdot)\|_{C_{c}^{a}(\mathbb{R}^{N})}\leq C_{1}(t_{0})\|f\|_{\mathrm{L}^{2}(\Omega)},$
  • ((b))
    for each $t_{0}>0$
@@
for any $\epsilon\in(0,a)$. Here $C_{1}(t_{0})$$C_{1,j}(t_{0})$ are positive constants. Authors used various eigenfunction estimates corresponding to (6) [see [Gru15b], [Gru15a], [AGCV23]]. The following interior regularity results can be found in [FRRO16].
@@
Let $u_{f}\in\mathrm{L}^{2}\left((0,T);H^{a}(\Omega)\right)$ be the solution to (1) with $f\in\mathrm{L}^{2}(\Omega)$. Then the following holds.
@@
    For $t>0$, $u_{f}(t,\cdot)\in C^{a}(\mathbb{R}^{N})$ and, for every $\beta\in[a,1+2a)$, $u_{f}(t,\cdot)$ is of class $C^{\beta}(\Omega)$ and
    (10) $[u_{f}(t,\cdot)]_{C^{\beta}(\Omega_{\delta})}\leq C\delta^{a-\beta},\forall\delta\in(0,1).$
@@
    The function ${\frac{u_{f}(t,\cdot)}{d^{a}}}\Big|_{\Omega}$ can be continuously extended to $\overline{\Omega}$. Moreover, there exists $\alpha\in(0,1)$ such that ${\frac{u_{f}}{d^{a}}\in C^{\alpha}(\overline{\Omega})}$. In addition, for $\beta\in[\alpha,a+\alpha]$, the following estimate holds
    (11) $\left[\frac{u_{f}}{d^{a}}\right]_{C^{\beta}(\Omega_{\delta})}\leq C\delta^{a-\beta},\forall\delta\in(0,1).$
Here $\Omega_{\delta}:=\{x\in\Omega:d(x,\partial\Omega)\geq\delta\}$ and $\cdot]_{C^{\beta}(\Omega_{\delta})}$ denotes the Hölder seminorm:
$\left[\frac{u_{f}(t,x)}{d^{a}}\right]_{C^{\beta}(\Omega_{\delta})}=\sup\limits_{x,y\in C^{\beta}(\Omega_{\delta}),x\neq y}\frac{|D^{\lfloor\beta\rfloor}u_{f}(t,x)-D^{\lfloor\beta\rfloor}u_{f}(t,y)|}{|x-y|^{\beta-\lfloor\beta\rfloor}}$
@@
(12) $\int_{\Omega}(x\cdot\nabla u_{f})(-\Delta)^{a}u_{f}\,\rm{d}x=\frac{2a-N}{2}\int_{\Omega}$ $u_{f}(-\Delta)^{a}u_{f}\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{u_{f}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma,$
@@ 1.3. $\mu$ transmission spaces
All the notations and definitions here are well explained in [Gru14], [Gru15a], [Gru16b], [Gru16a], [Gru18a], [Gru20]. We only recall some important definitions and concepts here.
We consider operators acting on the upper or lower half subspaces of $\mathbb{R}^{N}$ denoted by ±N:={(x1,,xN)N:xN
0
@@
. Similarly with respect to the boundary $\partial\Omega$ we divide a $C^{1,1}$ domain $\Omega$ into two parts. One part is the domain $\Omega$ itself and the other is the complement of it’s closure ${\mathbb{R}^{N}\setminus\overline{\Omega}}$. We denote $r^{\pm}$ as the restriction operator from $\mathbb{R}^{N}$ to ${\mathbb{R}^{N}_{\pm}}$ ($\Big$(or from $\mathbb{R}^{N}$ to $\Omega$ respectively ${\mathbb{R}^{N}\setminus\overline{\Omega}}\Big)$. The extension by zero from ${\mathbb{R}^{N}\setminus\overline{\Omega}}$ to $\mathbb{R}^{N}$ ($\Big$(or from $\Omega$ respectively ${\mathbb{R}^{N}\setminus\overline{\Omega}}$ to $\mathbb{R}^{N}\Big)$ is denoted by $e^{\pm}$.
@@
(13) $\mathcal{E}_{\mu}\left(\overline{\Omega}\right):=e^{+}\left\{u(x)=\tilde{d}^{\mu}v(x):v\in C^{\infty}\left(\overline{\Omega}\right)\right\},$
[…]
@@
$S^{m}_{1,0}:=\left\{\begin{aligned} p(x,\xi)\in C^{\infty}\left(\mathbb{R}^{N}\times\mathbb{R}^{N}\right):&\ |\partial_{x}^{\beta}\partial_{x}^{\alpha}p(x,\xi)|\sim O\left(\langle\xi\rangle^{m-|\alpha|}\right),\\ &\forall\alpha,\beta\in\mathbb{N}\cup\{0\},\ \mbox{for some}\ m\in\mathbb{C}\end{aligned}\right\}$
[…]
@@
. Hence by Paley-Wiener theorem ${\Xi^{t}_{\pm}}$ preserves support in ${\mathbb{R}^{N}_{\pm}}$. Similarly one can defined ${\Lambda^{t}_{\pm}}$ adapted to the situation for a smooth domain. For details readers are refereed to [Hor63], [Hör66], [Tay06], [TT91] [Gru08]. Here we mention some important homeomorphism results regarding this operators: for all $s>0$
@@
The space ${{\mathcal{E}_{\mu}\left(\overline{\Omega}\right)}}$ is dense in ${H^{\mu(s)}\left(\overline{\Omega}\right)}$ i.e., [see [Hor63], [Hör66], [Gru08]]
@@
Some further characterization can be found in [Gru15a], [Gru14]. These conditions are stated in the lemma below:
Lemma 1.
The following properties hold:
  • a:
    For all $s>\mu-\frac{1}{2}$, $H^{\mu(s)}(\overline{\Omega})\subset\overline{H}^{\mu-\frac{1}{2}}(\Omega)$ with continuous inclusion.
  • b:
    $H^{\mu(s)}(\overline{\Omega})=\dot{H}^{\mu+\frac{1}{2}-\epsilon}(\overline{\Omega})$, for all $s>\mu-\frac{1}{2}$ and $\epsilon>0$.
  • c:
    $H^{\mu(s)}(\overline{\Omega})=\dot{H}^{s}(\overline{\Omega})$ if ${\mu\in\left(s-\frac{1}{2},s+\frac{1}{2}\right)}$.
  • d:
    Let, $q\in C_{c}^{\infty}(\Omega)$. The operator ${r^{+}\big((-\Delta)^{s}+q\big)}$ is a homeomorphism from $H^{\mu(s)}(\overline{\Omega})$ onto $H^{\mu-2s}(\Omega)$.
  • e:
    $\overline{H}^{s}(\Omega)\subset H^{\mu(s)}(\overline{\Omega})\subset H^{s}_{loc}(\Omega)$ with continuous inclusions; i.e., multiplication by any $\chi\in C_{c}^{\infty}(\Omega)$ is bounded $H^{\mu(s)}(\overline{\Omega})\to H^{s}(\Omega)$.
The domain of fractional Laplacian with homogeneous Dirichlet boundary condition can be characterize as follows:
Lemma 2 ([CW20], [Gru18b]).
Let $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) be a bounded open set with $C^{1,1}$ boundary and $a\in(0,1)$. Let $D((-\Delta)^{a})$ be the domain of the operator $(-\Delta)^{a}$ with homogeneous Dirichlet boundary condition outside $\Omega$. Then, the following assertions holds.
  • (a)
    If $a\in(0,1/2)$, then $D((-\Delta)^{s})=\dot{H}^{2a}(\Omega)$.
  • (b)
    If $a=1/2$, then $D((-\Delta)^{a})\subset H^{1-\varepsilon}(\Omega)\cap\dot{H}^{a}(\Omega)$ for all $\varepsilon\in(0,1)$.
  • (c)
    If $a\in(1/2,1)$, then $D((-\Delta)^{a})\subset H^{a+\frac{1}{2}}(\Omega)\cap\dot{H}^{a}(\Omega)$.
@@
Let $f\in\mathrm{L}^{2}(\Omega)$, then the unique solution $u_{f}$ to (1) has the following regularity:
$u_{f}\in C\left((0,T);H^{a(2a)}(\overline{\Omega})\right)\cap C^{1}((0,T);\mathrm{L}^{2}(\Omega))\cap C([0,T);\mathrm{L}^{2}(\Omega)).$
It follows from semigroup theory [Paz12], [BB11].
@@
The theorem 7 is also true for the solution of the adjoint system corresponding to (1), given by:
(14) $\left\{\begin{aligned} -\partial_{t}u_{f^{*}}+(-\Delta)^{a}u_{f^{*}}+qu_{f^{*}}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{f^{*}}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\Omega\\ u_{f^{*}}(T,\cdot)=&f^{*}\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right$.
where $f^{*}\in\mathrm{L}^{2}(\Omega)$. It just comes from change of variable $t\to T-t$. We like to emphasize that Pohozaev identity (12) holds for $u_{f^{*}}$ also. Next we move onto defining traces of the transmission spaces.
1.4. Boundary values of $\mu$ transmission spaces, Dirichlet trace
The definitions and notations used in this sub-section are well explained in [Gru15a]. The author in [Gru15a], explained that the $\mu$ transmission spaces $H^{\mu(s)}$ admits a special definition of $\mu$ boundary values, which characterize how the elements of the $\mu$ transmission spaces decays towards the boundary. Let $M$ be a positive integer. We recall the space $\mathcal{E}_{\mu}$ as defined in (13). Let us introduce the natural canonical mapping
(15) $\rho_{\mu,M}:\mathcal{E}_{\mu}\to\mathcal{E}_{\mu}/\mathcal{E}_{\mu+M},$
where one can represent the space ${\frac{\mathcal{E}_{\mu}}{\mathcal{E}_{\mu+M}}}$ as the space of sections of a trivial bundle and introduce norms in it. Let us choose a smooth function $\tilde{d}$ on $\overline{\Omega}$ which is equal to the distance from the boundary sufficiently close to the boundary and remains positive in $\Omega$. We set
$I^{\mu}(x):=\frac{\tilde{d}^{\mu}(x)}{\Gamma(\mu+1)}\ \mbox{in}\ \overline{\Omega}\ \ \mbox{and}\ I^{\mu}\equiv 0\ \mbox{in}\partial\Omega,$
[…]
$u=u_{0}I^{\mu}+u_{1}I^{\mu+1}+\cdots+u_{M-1}I^{\mu+M-1}+v,$
where the functions $u_{i}\in C^{\infty}(\overline{\Omega})$ and constant along the normal closed to the boundary $\partial\Omega$. So, the traces of $u$ denoted by the following notation
$\gamma_{\mu,j}u=u_{j}|_{\partial\Omega}.$
Note that the fact that closed to the boundary, $I^{\mu}$ is equal to ${\frac{d^{\mu}}{\Gamma(\mu+1)}}$. Furthermore, note that:
$\gamma_{\mu,j}u=$ $\gamma_{\mu+j,0}u,\quad\mbox{when}\ u\in\mathcal{E}_{\mu+j},$
$\gamma_{\mu,0}u=$ $\Gamma(\mu+1)\gamma_{0}d^{-\mu}u,\quad\mbox{when}\ u\in\mathcal{E}_{\mu},\ \mbox{with}\ Re(\mu)>-1.$
where $\gamma_{0}$ is the projection of a function on the boundary $\partial\Omega$. We can extend the boundary trace for the functions in $H^{\mu(s)}(\overline{\Omega})$. For a function $u\in H^{\mu(s)}(\overline{\Omega})$, the boundary term $\frac{u}{d^{\mu}}$ is well defined and lies in the space $H^{s-\mu-\frac{1}{2}}(\partial\Omega)$ [see [Gru17], [Gru18a]]. The boundary trace $\frac{u}{d^{\mu}}$ for the function $u\in H^{\mu(s)}(\overline{\Omega})$, we call it as Dirichlet trace. The following result can be found in [Gru15a]:
Theorem 9 ([Gru15a]).
[…]
$\rho_{\mu,M}:H^{\mu(s)}(\overline{\Omega})\to\prod_{0\leq j<M}H^{s-Re(\mu)-j-\frac{1}{2}}(\partial\Omega).$
Furthermore, the map is surjective and with kernel $H^{(\mu+M)(s)}(\overline{\Omega})$. In other words, the canonical map $\rho_{\mu,M}$ defines a homeomorphism of ${H^{\mu(s)}(\overline{\Omega})/H^{(\mu+M)(s)}(\overline{\Omega})}$ onto ${\prod_{0\leq j<M}H^{s-Re(\mu)-j-\frac{1}{2}}(\partial\Omega)}$.
The theorem implies that if a function $u\in H^{\mu(s)}(\overline{\Omega})$ has zero Dirichlet trace, with $s>Re(\mu)+M-\frac{1}{2}$, then $u\in H^{(\mu+1)(s)}(\overline{\Omega})$. Similarly, we can define the Neumann trace also for the transmission spaces $H^{(a-1)(2a)}(\overline{\Omega})$.